Hemophilia is an x-linked recessive trait.  In a population of 2000 people, 4 women have hemophilia, 30 men have hemophilia, and 80 are carriers of the disorder. What is the frequency of the hemophilia allele? (Assume the population is 50% men and 50% women.)

a.

0.038

b.

0.059

c.

0.033

d.

0.057

e.

0.039
Answer #1

Answer:

e. 0.039

Explanation:

4 hemophilic women = 8 hemophilic alleles

30 hemophilic men = 30 hemophilic alleles

80 carrieres = 80 normal alleles and 80 hemophilic alleles

Remaining individuals are normal = 2000 – 114 = 1886

943 men = 943 normal alleles

943 women = 943*2 = 1886 normal alleles

Total alleles = 8+30+80+80+943+1886=3027

Total hemophilic alleles = 8+30+80 = 118

Frequency of hemophilic allele = 118/3027 = 0.039